40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15420     Entry time: Fri Jun 19 19:21:25 2020
Author: gautam 
Type: Update 
Category: General 
Subject: PSL shutter re-opened 

The PSL shutter was closed from the vacuum interlock trip. Today, I did the following:

  • Re-aligned input beam to PMC to recover high transmission / low reflection.
  • Re-set the LSC offsets.
  • ETMX watchdog was tripped. Reset it.
  • Opened the PSL shutter, IMC autolocker was able to lock the cavity almost immediately.
  • Tested POX/POY locking, ran the ASS to maximize single arm transmission.

All looks good for now. I will probably get back to PRFPMI locking Monday.

ELOG V3.1.3-